# Specific heat and density anomaly in the Hubbard model

### v1

@@ Abstract
Understanding thermal properties of materials is fundamental to technological applications and to discovering new phenomena. In particular, advances in experimental techniques such as cold-atom measurements allow the simulation of paradigmatic Hamiltonians with great control over model parameters, such as the Fermi-Hubbard model. One aspect of this model, which is not much explored, is the behavior of the specific heat as a function of density. In this work, we perform Determinant Quantum Monte Carlo simulations of the Hubbard model to analyze the specific heat as the filling, interaction, and temperature of the system are changed. We found that, with strong correlations, the specific heat presents a three-maxima structure as a function of filling, with local minima between them. This effect can be explained by a decomposition of kinetic and potential contributions to the specific heat, demonstrating interesting phenomena away from the commonly studied half-filling regime. Moreover, by analyzing the kinetic contribution in momentum space we show that, connected to this specific heat behavior, there is a density anomaly detected through the thermal expansion coefficient. These momentum-space quantities are accessible using cold-atom experiments. Finally, we map the location of these phenomena and connect the thermal expansion anomaly with the well-known Seebeck coefficient change of sign. Our results provide a new perspective to analyze this change of sign.
@@ I Introduction
Anomalous thermal properties of materials have become an intense source of research in different realms of physics. In the context of complex fluids, water presents a plethora of thermodynamic anomalies, including its unusual volume contraction from 0C to 4C at 1 atm [1], named as density anomaly, and its high specific heat. Both, among others, are regarded as essential ingredients for life [2]. Another paradigmatic example is the unexpected low thermal expansion of Fe–Ni Invar [3], a product of the competition between phonons and spins, which still pushes forward theoretical research and the development of materials where extreme precision is required [4]. Heavy fermion compounds, such as CeRu2Si2 [5], also display a negative electronic thermal expansion as well as an unusually high Grüneiser parameter, a quantity proportional to the ratio of thermal expansion to specific heat [6]. Organic charge-transfer salts [7, 8], like the spin liquid candidate $\kappa$-(BEDT-TTF)2Cu2(CN)3, exhibit anomalous expansivities depending on directionality, associated with a peak anomaly around 6 K which is accompanied by a hump in the specific heat. The electronic ingredient to the thermal expansion of such families of organic charge-transfer salts can be modeled with an effective Hubbard Hamiltonian [9] on an anisotropic triangular lattice at half-filling, where its parameters are connected to lattice constants.
@@
In this research paper, we report on simulations of the two-dimensional Fermi-Hubbard model in a square lattice, presenting results for all lattice fillings and a wide range of interactions. We demonstrate that: i) as correlations between electrons increase, a three-peak structure develops in the specific heat as a function of lattice filling, at $n=0.5$, $n=1$, and $n=1.5$. The peak in the central density is sharp, in connection with the known behavior as a function of temperature, while the lateral peaks are broader. ii) For strong correlations, there is a “density anomaly”, marked by a sign change of the thermal expansion coefficient $\alpha_{\mu}$ as doping varies. We address the connection between these two effects through an analysis of the momentum distribution of the particles. iii) Finally, we establish a mechanism associating the described thermal expansion anomaly with peaks in the entropy as a function of density $n$ and also the Seebeck anomaly, a relevant transport probe recently studied in this model [22, 23, 24, 25]. This quantity is known to exhibit changes of sign as a function of filling $n$ and temperature $T$. Moreover, it increases significantly close to half-filling. The analysis of $\alpha_{\mu}$ sheds light on both of these characteristic behaviors.
@@ II Model and Methods
where $i$ runs over all lattice sites and $\braket{i,j}$ correspond to first neighbor sites. Operators $c^{\dagger}_{\bm{i}\sigma}$/$c_{\bm{i}\sigma}$ denote the creation/annihilation of electrons of spin $\sigma$ on a site $\bm{i}$, and the number operator is $n_{\bm{i}\sigma}\equiv c^{\dagger}_{\bm{i}\sigma}c_{\bm{i}\sigma}$. The first sum describes particle hopping processes with amplitude $t$, the second term controls the density through the chemical potential $\mu$, and the last term describes an on-site repulsive interaction $U$. The Hamiltonian of Equation 1 in bipartite lattices, such as the square lattice, has particle-hole symmetry. That is, $n(\mu)=2-n(-\mu)$. At $\mu=0$ the system is at half-filling, $n(T,U,\mu=0)=1$, and its ground state is a Mott insulator. Due to this symmetry, we mostly focus on the region $n\leq 1$. As $U$ hinders doubly occupied sites, a chemical potential $\mu\sim U$ is required to add another particle to the system [27, 28, 29, 30], leading to the formation of a Mott plateau around half-filling.
Our simulations were performed using the Determinant Quantum Monte Carlo (DQMC) method [31, 32, 33, 34, 35, 36], an unbiased technique. It uses a discrete auxiliary field to break down the interaction term of the $d$-dimension many-body problem into a non-interacting $(d+1)$-dimension Hamiltonian with $L$ imaginary time slices, introducing an error of $\mathcal{O}(\Delta\tau^{2})$, where $\beta=L\Delta\tau$ is the inverse temperature. Our results are obtained from 10 independent simulations for each parameter set, each simulation with 2000 equilibration steps and up to 50000 measurement steps, choosing $\Delta\tau\leq 0.1$ to ensure the Trotter error is smaller than the statistical error from the Monte Carlo sampling. The square lattice size is $10\times 10$ to calculate thermodynamic quantities. We set $t=1$ as our energy scale. The lattice constant is also set to unity. To determine the specific heat, we numerically differentiate the internal energy at constant density using finite differences:
@@
where $c_{\mu}=\left(\frac{\partial E}{\partial T}\right)_{\mu}$ is the specific heat at constant chemical potential, $\alpha_{\mu}=\left(\frac{\partial n}{\partial T}\right)_{\mu}$ is the thermal expansion coefficient, $\kappa_{T}=\frac{1}{n^{2}}\frac{dn}{d\mu}$ is the isothermal compressibility, $E(T,\mu)$ is the internal energy of the system, and $n\equiv\braket{n}$ is the average number of particles per lattice site. In particular, for $T=0.5$ we used $\Delta\tau=0.05$, $L=36,40,44$.
@@
Using finite-differences for different temperatures, we obtain $\left(\frac{\partial n_{\bm{k}}}{\partial T}\right)_{\mu}$. To obtain a better resolution when plotting quantities in momentum space, this is calculated for a $24^{2}$ lattice.
@@ III Results
The specific heat $c_{n}$ as a function of doping $n$ is exhibited in Figure 1 a). In the weak-coupling regime, it has a maximum at $n=1$. As the interaction $U$ becomes stronger and beyond $U\approx 6$, a three-peak structure emerges: two lateral peaks around $n=0.5$ and $n=1.5$, and a central maximum at half-filling. Enhancing the on-site repulsion makes the central peak sharper, and the local minima around it move towards half-filling. In both analytically solvable scenarios, $U=0$ (non-interacting case) and $t=0$ (atomic limit), such a triple-peak structure is absent. Therefore, the reported behavior emerges from the electronic correlations and the interplay between $U$ and $t$. In experimental measurements, specific heat peaks at finite doping have been detected for some materials [19].
@@
Figure 1: a) Specific heat at constant density $c_{n}$ as a function of lattice filling for $T=1$ and various values of interaction $U$. b) and c) are the kinetic and potential energy contributions to the specific heat for the same temperature and interaction strengths. Error bars are smaller than the symbol size.
@@
Figure 2: Specific heat and kinetic/potential contributions for $U=10$ at temperatures a) $T=2$, b) $T=1$ and c) $T=0.5$. The inset in b) is a $T\times U$ ($0\leq U\leq 10$, $0\leq T\leq 2.5$) plot at half-filling with dashed lines showing the temperatures for which we display specific results in a), b), and c). Red circles and purple squares represent, respectively, the high and low $T$ peaks in the specific heat as a function of temperature, with data as a function of temperature extracted from [17].
@@
The system’s average energy $E$ per lattice site can be written as the sum of the kinetic part $E_{K}=\sum_{\bm{k}}\varepsilon_{\bm{k}}n_{\bm{k}}$, where $\varepsilon_{\bm{k}}$ is the lattice dispersion relation and $n_{\bm{k}}$ is the momentum distribution of particles, and the potential term $E_{P}=Ud$, where $d$ is the average double occupancy $d=\braket{n_{\uparrow}n_{\downarrow}}$. Following this idea, we decompose the specific heat into its kinetic and potential energy contributions, as shown in Figure 1 b) and c) respectively. The kinetic energy portion $\left(\frac{\partial E_{\text{K}}}{\partial T}\right)_{n}$ is nearly independent of the interaction for low densities $0\leq n\leq 0.5$. Thus, this regime resembles the non-interacting scenario, where the electrons are too spread out for the $U$ term to be relevant. For $10.5<n<$ 1, the effects of the repulsion become important (double occupancy starts to build up), and the kinetic energy contribution starts to decrease as the interaction increases. Hence, for large $U$ we have a maximum around $n=0.5$ (and $n=1.5$, by symmetry), whereas for weak interactions there is a single maximum at half-filling.
@@
In Figure 2, we plot the total $c_{n}$ along with its kinetic and potential contributions for $U=10$, considering three different temperatures: $T=2$, $T=1$, and $T=0.5$. Concerning the higher temperature $T=2$, $c_{n}$ displays only a central peak illustrated in Figure 2 a) for this particular repulsion value. As the temperature gets lower, lateral peaks at $n=0.5$ and $n=1.5$ start to develop according to Figure 2 b), and their magnitude gets enhanced at even lower temperatures as shown in Figure 2 c). At such low temperatures, the potential energy contribution to $c_{n}$ is negative for all fillings, even at $n=1$. This implies that $\left(\frac{\partial d}{\partial T}\right)_{n}<0$, an effect particularly important in cold atoms experiments exploring adiabatic cooling: whereas $\partial d/\partial U=-\partial d/\partial T$, a change in $U$ can be used to decrease the temperature at constant entropy. At half-filling, this has been studied for square [37, 29], kagome [38], and honeycomb [39, 40] lattices, but the results of panels b) and c) indicate that a small doping can enhance the effect. The inset in Figure 2 b) is a $T\times U$ plot at half-filling with dashed lines showing the temperatures for which we display results. Red circles and purple squares represent, respectively, the high and low $T$ peaks in the specific heat as a function of temperature. Hence, at our higher temperature, $T=2$, we cross the high $T$ peak for large $U$. At our intermediate temperature $T=1$, we cross the high $T$ peak for small $U$. And our lower temperature, $T=0.5$, does not cross any peak as a function of temperature at half-filling. In short, the decomposition highlights the different nature of the central and lateral peaks. The lateral, broader ones come from the strong decrease in the kinetic energy due to reduced carrier mobility. The central sharp peak is due to an increase in double occupancy for small to intermediate temperatures, and decreases as $T$ decreases due to the formation of local moments.
@@
We analyze finite-size effects of the specific heat in Figure 3. We fixed $U=10$ and $T=1$, and considered lattice sizes from $6^{2}$ to $24^{2}$. Results coincide within error bars, indicating that our thermodynamic analysis in $10^{2}$ lattices provides a good estimate of the thermodynamic limit results. We note that even the half-filling peak remains the same for large sizes, although it could be expected to increase in the thermodynamic limit, for lower temperatures than considered here, due to the known results at $n=1$ [17]. In our results, this does not happen because the temperature $T=0.5$ sits between the peaks as a function of temperature, as shown in the inset of Figure 2 b).
@@
Figure 3: Specific heat as a function of lattice filling $n$ for $U=10$, $T=1$ and different lattice sizes.
Inasmuch as the decrease of $c_{n}$ in the region $10.5<n<$ 1 is dominated by the kinetic contribution and
@@
we analyze in Figure 4 the temperature derivative of $n_{\bm{k}}$ within the first Brillouin zone. The six panels cover weak and strong couplings as the density varies. At a fixed filling, large $U$ increases correlations in the real domain and delocalizes electrons in momentum space. Consequently, correlations induce a damping in the magnitude of $\left(\frac{\partial n_{\bm{k}}}{\partial T}\right)_{n}$. This effect gets more pronounced as one approaches the Mott state, where the system gets fully localized in real space. This is more clearly seen in the absolute value of the gradient, shown in Figure 5. The lower sensitivity to temperature at high density and strong coupling is reflected in the kinetic energy derivative with temperature of Equation 4, which is lower for $n=0.83$ than for $n=0.51$. This explains the reduction of $c_{n}$ beyond the quarter filling. Commonly addressed in cold atoms experiments, the momentum distribution [41] imaging is accessible through time-of-flight measurements [42, 43, 44]. Double occupancy measurements for fermionic atoms with atomic resolution are available through quantum gas microscopy [45, 46, 47, 48].
@@
We note that we plotted $\left(\partial n_{\bm{k}}/\partial T\right)_{\mu}$, but the kinetic energy contribution would actually be calculated from $\left(\partial n_{\bm{k}}/\partial T\right)_{n}$. We compared these two quantities (not shown), and they are qualitatively the same: strong interactions and higher densities lead to a decreased intensity in momentum space.
@@
at constant $\mu$. Figure 6 a) shows, for different lattice fillings and interaction strengths at a temperature $T=1$, a trivial crossing at $\alpha_{\mu}=0$ and $n=1$ due to particle-hole symmetry of the square lattice, since $n=1$ at $\mu=0$ for any temperature and interaction. For weak interactions (up to $U\approx 4$), we observe $\alpha_{\mu}>0$ below half-filling and $\alpha_{\mu}<0$ above it. This describes the regular response of the thermal expansion predicted by the non-interacting solution. Indeed, if $n<1$, raising the temperature increases the average occupation number at fixed chemical potential, while an increase in temperature tends to decrease the density when $n>1$. As $U$ becomes larger (beyond $U\approx 6$), the thermal expansion takes on negative values just below the half-filling and positive ones just above it. This composes the signature of the anomaly: $\alpha_{\mu}<0$ for $n<1$ and $\alpha_{\mu}>0$ for $n>1$. The reported anomalous sign change persists up to the atomic limit, which can be solved analytically. Figure 6 b) shows the density as a function of temperature at $|\mu|=1$. The red curves portray the non-interacting case $U=0$, where the particles’ thermal response is regular. The blue curves correspond to the intermediate coupling scenario $U=6$ for which the anomaly is observed at low temperatures $T\lesssim 2$. This kind of effect has been widely studied for classical liquid systems with short-range interactions at competing distance scales, most notably water, including on Monte Carlo simulations of classical lattice models [49, 50, 51]. In the context of Fermi-Hubbard physics, the anomalous thermal response arises from the strongly correlated particles’ motion near the Mott states [27, 28, 29, 30]. Recently, some of us proposed such an effect for a system of bosons [14, 10]. Thus, these DQMC results for the Hubbard model demonstrate that it can also happen for Fermi statistics.
Figure 6: a) Thermal expansion $\alpha_{\mu}$ as a function of density for $T=1$, considering interactions ranging from the weak to the strong coupling limit. b) Density as a function of temperature, showing a case with a regular density behavior (red circles, $U=0$) and an anomalous behavior (blue triangles, $U=6$). The continuous black line is the half-filling density. The vertical dotted line is $T=1$, corresponding to the fixed temperature in panel a).

### v2

@@ Abstract
Understanding thermal properties of materials is fundamental to technological applications and to discovering new phenomena. In particular, advances in experimental techniques such as cold-atom measurements allow the simulation of paradigmatic Hamiltonians with great control over model parameters, such as the Hubbard model. One aspect of this model which is not much explored is the behavior of the specific heat as a function of density. In this work, we perform Determinant Quantum Monte Carlo simulations of the Hubbard model interpolating between the square and triangular lattices to analyze the specific heat as the filling, interaction, and temperature of the system are changed. We found that, with strong correlations, the specific heat presents a three-maxima structure as a function of filling, with local minima between them. This effect can be explained by a decomposition of kinetic and potential contributions to the specific heat, demonstrating interesting phenomena away from the commonly studied half-filling regime. Moreover, by analyzing the kinetic contribution in momentum space we show that, connected to this specific heat behavior, there is a density anomaly detected through the thermal expansion coefficient. These momentum-space quantities are accessible using cold-atom experiments measurements at multiple temperatures. Finally, we map the location of these phenomena and connect the thermal expansion anomaly with the well-known Seebeck coefficient change of sign. Our results provide a new perspective to analyze this change of sign.
@@ I Introduction
Anomalous thermal properties of materials have become an intense source of research in different realms of physics. In the context of complex fluids, water presents a plethora of thermodynamic anomalies, including its unusual volume contraction from 0C to 4C at 1 atm [1], named as density anomaly, and its high specific heat. Both, among others, are regarded as essential ingredients for life [2]. Another paradigmatic example is the unexpected low thermal expansion of Fe–Ni Invar [3], a product of the competition between phonons and spins, which still pushes forward theoretical research and the development of materials where extreme precision is required [4]. Heavy fermion compounds, such as CeRu2Si2 [5], also display a negative electronic thermal expansion as well as an unusually high Grüneisen parameter, a quantity proportional to the ratio of thermal expansion to specific heat [6]. Organic charge-transfer salts [7, 8], like the spin liquid candidate $\kappa$-(BEDT-TTF)2Cu2(CN)3, exhibit anomalous expansivities depending on directionality, associated with a peak anomaly around 6 K which is accompanied by a hump in the specific heat. The electronic ingredient to the thermal expansion of such families of organic charge-transfer salts can be modeled with an effective Hubbard Hamiltonian [9] on an anisotropic triangular lattice at half-filling, where its parameters are connected to lattice constants.
@@
In this research paper, we report on simulations of the two-dimensional Hubbard model in a square lattice, presenting results for all lattice fillings and a wide range of interactions. We demonstrate that: i) as correlations between electrons increase, a three-peak structure develops in the specific heat as a function of lattice filling, at $n=0.5$, $n=1$, and $n=1.5$ in the square lattice. The peak in the central density is sharp, in connection with the known behavior as a function of temperature, while the lateral peaks are broader and are altered for different lattices. ii) For strong correlations, there is a “density anomaly”, marked by a sign change of the thermal expansion coefficient $\alpha_{\mu}$ as doping varies. We address the connection between these two effects through an analysis of the momentum distribution of the particles. iii) Finally, we clarify the mechanism associating the thermal expansion behavior with the Seebeck anomaly, a relevant transport probe recently studied in the Hubbard model [22, 23, 24, 25]. This quantity is known to exhibit changes of sign as a function of filling $n$ and temperature $T$. Moreover, it increases significantly close to half-filling. The analysis of $\alpha_{\mu}$ sheds light on both of these characteristic behaviors.
@@ II Model and Methods
where $i$ runs over all lattice sites and $\braket{i,j}$ correspond to first neighbor sites. Operators $c^{\dagger}_{\bm{i}\sigma}$$c_{\bm{i}\sigma}$ denote the creation/annihilation of electrons of spin $\sigma$ on a site $\bm{i}$, and the number operator is $n_{\bm{i}\sigma}\equiv c^{\dagger}_{\bm{i}\sigma}c_{\bm{i}\sigma}$. The first sum describes particle hopping processes with amplitude $t$, the second term controls the density through the chemical potential $\mu$, and the last term describes an on-site repulsive interaction $U$.
The Hamiltonian of Equation 1 in bipartite lattices, such as the square lattice, has particle-hole symmetry. That is, $n(\mu)=2-n(-\mu)$. At $\mu=0$ the system is at half-filling, $n(T,U,\mu=0)=1$, and its ground state is a Mott insulator. As $U$ hinders doubly occupied sites, a chemical potential $\mu\sim U$ is required to add another particle to the system, leading to the formation of a Mott plateau around half-filling [27, 28, 29, 30].
Our simulations were performed using the Determinant Quantum Monte Carlo (DQMC) method [31, 32, 33, 34, 35, 36], an unbiased technique. It uses a discrete auxiliary field to break down the interaction term of the $d$-dimensional many-body problem into a non-interacting $(d+1)$-dimensional Hamiltonian with $L$ imaginary time slices, introducing an error of $\mathcal{O}(\Delta\tau^{2})$, where $\beta=L\Delta\tau$ is the inverse temperature. Our results are obtained from 10 independent simulations for each parameter set, each simulation with 2000 equilibration steps and up to 50000 measurement steps, choosing $\Delta\tau\leq 0.1$ to ensure the Trotter error is smaller than the statistical error from the Monte Carlo sampling. The lattice size is $10\times 10$ to calculate thermodynamic quantities. We set $t=1$ as our energy scale. The lattice constant is also set to unity.
To determine the specific heat, we numerically differentiate the internal energy at constant density using finite differences:
@@
where $c_{\mu}=\left(\frac{\partial E}{\partial T}\right)_{\mu}$ is the specific heat at constant chemical potential, $\alpha_{\mu}=\left(\frac{\partial n}{\partial T}\right)_{\mu}$ is the thermal expansion coefficient, $\kappa_{T}=\frac{1}{n^{2}}\left(\frac{\partial n}{\partial\mu}\right)_{T}$ is the isothermal compressibility, $E(T,\mu)$ is the internal energy of the system, and $n\equiv\braket{n}$ is the average number of particles per lattice site (see Appendix A for the derivation of the equation).
@@
Using finite-differences for different temperatures, we obtain $\left(\frac{\partial n_{\bm{k}}}{\partial T}\right)_{\mu}$. To improve the resolution when plotting quantities in momentum space, this is calculated for a $24\times 24$ lattice.
@@ III Results
The specific heat $c_{n}$ as a function of doping $n$ is exhibited in Figure 1 a). In the weak-coupling regime, it has a maximum at $n=1$. As the interaction $U$ becomes stronger and beyond $U\approx 6$, a three-peak structure emerges: two lateral peaks around $n=0.5$ and $n=1.5$, and a central maximum at half-filling. Enhancing the on-site repulsion makes the central peak sharper, and the local minima around it move towards half-filling
In both analytically solvable scenarios, $U=0$ (non-interacting case) and $t=0$ (atomic limit), such a triple-peak structure is absent. Therefore, the reported behavior emerges from the electronic correlations and the interplay between $U$ and $t$. In experimental measurements, specific heat peaks at finite doping have been detected for some materials [19].
At half-filling, there is a well-known two-peak structure for $c_{n}(T)$ [17]: a spin peak at $T\sim J=4t^{2}/U$ and a charge peak $T\sim U$. Our simulations remain in temperatures above the spin scale. In particular, for our lowest $T=0.5$, we have $J\lesssim T\ll U$ and double occupancy is mostly suppressed. Hence, as $n$ is tuned away from half-filling, the main contribution to the energy should be from the hopping term, that is, the number of electron–hole bonds.
@@
Figure 1: a) Specific heat at constant density $c_{n}$ as a function of lattice filling for $T=1$ and various values of interaction $U$. b) and c) are the kinetic and potential energy contributions to the specific heat for the same temperature and interaction strengths. Results are for the $10\times 10$ lattice.
@@
Figure 2: Specific heat and kinetic/potential contributions for $U=10$ at temperatures a) $T=2$, b) $T=1$ and c) $T=0.5$. The inset in b) is a $T\times U$ ($0\leq U\leq 10$, $0\leq T\leq 2.5$) plot at half-filling with dashed lines showing the temperatures for which we display specific results in a), b), and c). Red circles and purple squares represent, respectively, the high and low $T$ peaks in the specific heat as a function of temperature, with data as a function of temperature of the inset extracted from reference [17]. Results are for the $10\times 10$ lattice.
@@
The system’s average energy $E$ per lattice site can be written as the sum of the kinetic part $E_{K}=\sum_{\bm{k}}\varepsilon_{\bm{k}}n_{\bm{k}}$, where $\varepsilon_{\bm{k}}$ is the lattice dispersion relation and $n_{\bm{k}}$ is the momentum distribution of particles, and the potential term $E_{P}=Ud$, where $d$ is the average double occupancy $d=\braket{n_{\uparrow}n_{\downarrow}}$. Following this idea, we decompose the specific heat into its kinetic and potential energy contributions, as shown in Figure 1 b) and c) respectively. The kinetic energy portion $\left(\frac{\partial E_{\text{K}}}{\partial T}\right)_{n}$ is nearly independent of the interaction for low densities $0\leq n\leq 0.5$. Thus, this regime resembles the non-interacting scenario, with electrons being too spread out for the $U$ term to be relevant. For $0.5<n<1$, the repulsion effects become important (double occupancy starts to build up), and the kinetic energy contribution starts to decrease as the interaction increases. Hence, for large $U$ we have a maximum around $n=0.5$ (and $n=1.5$, by symmetry), whereas for weak interactions there is a single maximum at half-filling.
@@
In Figure 2 we plot the total $c_{n}$ along with its kinetic and potential contributions for $U=10$, considering three different temperatures: a) $T=2$, b) $T=1$, and c) $T=0.5$. Concerning the higher temperature $T=2$, $c_{n}$ displays only a central peak, illustrated in Figure 2 a) for this particular repulsion value. This temperature is away from the $T\sim J$ region, leading to a decreased kinetic energy contribution and no lateral maxima in the total $c_{n}$. As the temperature reaches the region $J\lesssim T$, lateral peaks at $n=0.5$ and $n=1.5$ start to develop according to Figure 2 b). Their magnitude is slightly enhanced at even lower temperatures as shown in Figure 2 c), even though we are still at $J\lesssim T$.
At the lowest temperatures we considered, the potential energy contribution to $c_{n}$ is negative for all fillings. This implies that $\left(\frac{\partial d}{\partial T}\right)_{n}<0$, an effect particularly important in cold atoms experiments exploring adiabatic cooling: whereas $\partial d/\partial U=-\partial d/\partial T$, a change in $U$ can be used to decrease the temperature at constant entropy. At half-filling, this has been studied for square [37, 29], kagome [38], and honeycomb [39, 40] lattices. The results of panels b) and c) show that this can happen for the whole doping range.
This counterintuitive behavior is claimed to be analogous to the Pomeranchuk effect in liquid $^{3}\mathrm{He}$ [41]. For the Hubbard model, since the (spin) entropy is larger in a localized state than when the fermions form a Fermi liquid, it is favorable to increase the degree of localization upon heating. At low $T$, the system behaves as a coherent Fermi liquid with small entropy, while at higher temperatures, local moments increase with $T$, favoring electron localization, creating a Pomeranchuk-like tendency toward the higher-entropy local-moment regime.
The inset in Figure 2 b) is a $T\times U$ plot at half-filling with dashed lines showing the temperatures for which we display results. Red circles and purple squares represent, respectively, the high and low $T$ peaks in the specific heat as a function of temperature. Hence, at our higher temperature, $T=2$, we cross the high $T$ peak for large $U$. At our intermediate temperature $T=1$, we cross the high $T$ peak for small $U$. Our lower temperature, $T=0.5$, does not cross any peak as a function of temperature at half-filling, hence guaranteeing we are in the $J\lesssim T$ region for $T=0.5$ and $T=1$, while $T=2$ is above this spin regime for most interactions.
Therefore, the decomposition of $c_{n}$ for $J\lesssim T$ highlights the different nature of the central and lateral peaks. The lateral, broader ones come from the strong decrease in the kinetic energy due to reduced carrier mobility. The central sharp peak is due to an increase in double occupancy for small to intermediate temperatures, and decreases as $T$ decreases due to the formation of local moments. At $T\sim U$ only the central peak is detected, and at very low temperature, we expect all peaks to vanish.
@@
We analyze finite-size effects of the specific heat in Figure 3. We fixed $U=10$ and $T=1$ in panel a), and considered lattice sizes from $6\times 6$ to $24\times 24$. Results coincide within error bars, indicating that our analysis in $10\times 10$ lattices provides a good estimate of the thermodynamic limit results. For the lower temperature of $T=0.5$ shown in panel b) of Figure 3, the specific heat is also in good agreement within error bars. We note that even the half-filling peak remains the same for large sizes, although it could be expected to increase in the thermodynamic limit for lower temperatures than considered here, due to the known results at $n=1$ [17]. In our results, this does not happen because the temperature $T=0.5$ is situated between the peaks as a function of temperature, as shown in the inset of Figure 2 b).
@@
Figure 3: Specific heat as a function of lattice filling $n$ for $U=10$ and different lattice sizes at a) $T=1.0$ and b) $T=0.5$.
Inasmuch as the decrease of $c_{n}$ in the region $0.5<n<1$ is dominated by the kinetic contribution and
@@
we analyze in Figure 4 the temperature derivative of $n_{\bm{k}}$ within the first Brillouin zone. The six panels cover weak and strong couplings as the density varies. At a fixed filling, large $U$ increases correlations in the real domain and delocalizes electrons in momentum space. Consequently, correlations induce a damping in the magnitude of $\left(\frac{\partial n_{\bm{k}}}{\partial T}\right)_{n}$. This effect gets more pronounced as one approaches the Mott state, where the system gets fully localized in real space. This is more clearly seen in the absolute value of the gradient, shown in Figure 5.
@@
The lower sensitivity to temperature at high density and strong coupling is reflected in the kinetic energy derivative with temperature of Equation 4, which is lower for $n=0.83$ than for $n=0.51$. This explains the reduction of $c_{n}$ beyond the quarter filling. Commonly addressed in cold atoms experiments, the momentum distribution [42] imaging is accessible through time-of-flight measurements [43, 44, 45], but measurements at multiple temperatures [46] would have to be performed to calculate the temperature derivative $\partial n\bm{k}/\partial T$. Double occupancy measurements for fermionic atoms with atomic resolution are available through quantum gas microscopy [47, 48, 49, 50], again requiring measurements at different temperatures for a numerical differentiation.
@@
While Figures 4 and 5 highlight the reduced temperature sensibility for strong interactions at $n\approx 0.8$, we need $\varepsilon(\bm{k})\left(\frac{\partial n_{\bm{k}}}{\partial T}\right)_{n}$ to understand which regions contribute to the specific heat (see Appendix D for a numerical comparison of $\left(\partial n_{\bm{k}}/\partial T\right)_{\mu}$ and $\left(\partial n_{\bm{k}}/\partial T\right)_{n}$, they are qualitatively the same). We plot this in the left panels of Fig. 6, for $n\approx 0.83$ at $T=1$. For $U=2$, the largest contribution comes from small $k$s. Increasing the interaction to $U=10$ weakens this central feature, while the border regions remain mostly unaltered. Hence, the lower specific heat is due to a decrease in the thermal sensitivity of low $k$ electrons at the center of the Brillouin zone.
Note that the entropy is related to $\left(\frac{\partial n}{\partial T}\right)_{\mu}$:
$s(T,\mu)=\int_{-\infty}^{\mu}\left(\frac{\partial n}{\partial T}\right)_{\mu^{\prime}}d\mu^{\prime}$ (5)
Since $n=\sum_{\bm{k}}n_{\bm{k}}$, we can decompose it as
$s(T,\mu)$ $=\sum_{\bm{k}}\int_{-\infty}^{\mu}\left(\frac{\partial n_{\bm{k}}}{\partial T}\right)_{\mu^{\prime}}d\mu^{\prime}$ (6)
$=\sum_{\bm{k}}s_{\bm{k}}(T,\mu)$ (7)
In the right panels of Fig. 6 we show the momentum space decomposition of the entropy, $s_{\bm{k}}$ at a density $n\approx 0.83$ and temperature $T=1$ for the weakly interacting ($U=2$ and strong coupling ($U=10$) regimes. Due to the $\mu^{\prime}$ integration required, we are only able to perform this calculation for $10\times 10$ systems. Surprisingly, at strong coupling, we see a negative contribution to the entropy from the electrons at the center of the Brillouin zone, while the positive entropy contributions come from the Fermi surface region. This is consistent with the decreasing entropy as a function of $n$ observed in other works [28, 22] and shown for various temperatures in Appendix C.
Figure 6: Left panels show the momentum space contributions to $\varepsilon(\bm{k})\left(\frac{\partial n_{\bm{k}}}{\partial T}\right)_{n}$. Right panels shows the momentum space contributions to the entropy. The interaction is $U=2$ (upper panels) and $U=10$ (lower panels), all of them for a filling $n\approx 0.83$. The temperature is $T=1$ and the lattice size is $24\times 24$ on the left panels, with a cubic spline interpolation between points. The right panels are calculated for a $10\times 10$ lattice due to $\mu$ integration needed to find $s$.
@@
at constant $\mu$. Figure 7 a) shows, for different lattice fillings and interaction strengths at a temperature $T=1$, a trivial crossing at $\alpha_{\mu}=0$ and $n=1$ due to particle-hole symmetry of the square lattice, since $n=1$ at $\mu=0$ for any temperature and interaction. For weak interactions (up to $U\approx 4$), we observe $\alpha_{\mu}>0$ below half-filling and $\alpha_{\mu}<0$ above it. This describes the regular response of the thermal expansion predicted by the non-interacting solution. Indeed, if $n<1$, raising the temperature increases the average occupation number at fixed chemical potential, whereas increasing the temperature tends to decrease the density when $n>1$. As $U$ becomes larger (beyond $U\approx 6$), the thermal expansion takes on negative values just below the half-filling and positive ones just above it. This composes the signature of the anomaly: $\alpha_{\mu}<0$ for $n<1$ and $\alpha_{\mu}>0$ for $n>1$. The reported anomalous sign change persists up to the atomic limit, which can be solved analytically.
Figure 7 b) shows the density as a function of temperature at $|\mu|=1$. The red curves portray the non-interacting case $U=0$, where the particle’s thermal response is regular. The blue curves correspond to the intermediate coupling scenario $U=6$ for which the anomaly is observed at low temperatures $T\lesssim 2$. This kind of effect has been widely studied for classical liquid systems with short-range interactions at competing distance scales, most notably water, and on Monte Carlo simulations of classical lattice models [51, 52, 53]. In the context of Hubbard physics, the anomalous thermal response arises from the strongly correlated particles’ motion near the Mott states [27, 28, 29, 30]. Recently, some of us proposed such an effect for a system of bosons [14, 10]. Thus, these DQMC results for the Hubbard model demonstrate that it can also happen for Fermi statistics.
Figure 7: a) Thermal expansion $\alpha_{\mu}$ as a function of density for $T=1$, considering interactions ranging from the weak to the strong coupling limit. b) Density as a function of temperature, showing a case with a regular density behavior (red circles, $U=0$) and an anomalous behavior (blue triangles, $U=6$). The continuous black line is the half-filling density. The vertical dotted line is $T=1$, corresponding to the fixed temperature in panel a). Results are for the $10\times 10$ lattice.
@@
In particular, $S_{\text{Kelvin}}$ arises from the slow limit of the general Kubo formula for thermopower, that is, $\lim\{q_{x}\rightarrow 0,\omega\rightarrow 0\}\;S(\bm{q},\omega)$. In the context of the Hubbard model, it has been established that this approximation is one of the best available [54], even being in excellent quantitative agreement with experiments for some materials [56].